# Optimal Solutions for the Moving Target Vehicle Routing Problem with Obstacles via Lazy Branch and Price

### 23 Mar 2026 v1

@@ I INTRODUCTION
Traditionally, the selection step within column generation (known as the restricted master problem) aims to minimize the sum of selected tours’ costs. In the MT-VRP-O, however, computing tour costs is expensive, since it requires collision-free motion planning. Our key idea is to instead perform column generation using cheap-to-compute lower bounds on tour costs. We compute these lower bounds by solving a motion planning problem with relaxed continuity constraints. Thus, we incorporate an outer alternation between (i) column generation using lower bounds on tour costs, and (ii) lazily evaluating only the costs of tours selected by column generation. We evaluate the cost of a tour by searching for a shortest path in a Graph of Convex Sets (GCS) [12]; we use our continuity relaxation strategy to provide a heuristic for the search. If column generation selects a set of tours whose costs have all been evaluated, we terminate the alternation between (i) and (ii). Our numerical results show that Lazy BPRC runs up to 44 times faster than a non-lazy ablation, and up to 13 times faster than an ablation using an existing obstacle-unaware heuristic [13].
@@ V-D Computing Lower Bound on Tour Cost
This section describes how we compute the lower bound $\underline{c}(\Gamma)$ for a tour $\Gamma$ in Section V. Our method extends the procedure from BPRC [8] to handle obstacles, and we illustrate our method in Fig. 2. At the beginning of Lazy BPRC, we divide each target-window $\gamma_{i,j}$ into segments, where $\xi_{i,j,k}=(i,[\underline{t}_{i,j,k},\overline{t}_{i,j,k}])$ denotes the $k$th segment of $\gamma_{i,j}$, and ${\textnormal{Segments}}(\gamma_{i,j})$ is the set of segments of $\gamma_{i,j}$. To determine the number of segments per target-window, we first specify a number of segments to allocate per target, denoted as $n_{\text{seg,tar}}$. Then for each target $i$, we allocate segments among its windows using the formula from BPRC [8], which allocates more segments to longer windows. The depot gets a single segment $\xi_{0}$. For a target-window $\gamma$, $n_{\text{seg}}(\gamma)$ denotes the number of segments allocated to $\gamma$. The segment indices for a target-window are ordered in increasing order of start time. We say an agent trajectory $\tau_{\textnormal{a}}$} intercepts segment $\xi_{i,j,k}$ if $\tau_{\textnormal{a}}$} intercepts target $i$ at a time $t\in[\underline{t}_{i,j,k},\overline{t}_{i,j,k}]$.
@@
Figure 2: Computing bounds on the cost of an example tour $\Gamma$. To compute the lower bound $\underline{c}(\Gamma)$, we divide each target-window visited by $\Gamma$ into segments. We then construct a segment-graph $\mathcal{G}_{\text{seg}}$, where the nodes are the segments, and an edge connects every segment of $\Gamma[n]$ to every segment of $\Gamma[n+1]$. The edge cost from segment $\xi$ to $\xi^{\prime}$ is the distance traveled along the shortest path in space from $\xi$ to $\xi^{\prime}$, if this path satisfies the relaxed timing constraints in Section V-D, and $\infty$ otherwise. $\underline{c}(\Gamma)$ is the cost of the shortest path in $\mathcal{G}_{\text{seg}}$ from $\xi_{0}$ to $\xi_{0}$ visiting all target-windows in $\Gamma$. To compute the upper bound $\overline{c}$, we construct a segment-start-graph $\mathcal{G}_{\text{start}}$, where the nodes are the starting points of the segments, and an edge connects every segment-start of $\Gamma[n]$ to every segment-start of $\Gamma[n+1]$. The edge cost from $s$ to $s^{\prime}$ is distance traveled by a feasible minimum-distance agent trajectory from $s$ to $s^{\prime}$, if such a trajectory exists or if $s^{\prime}=s_{0}$, and $\infty$ otherwise. Our upper bound is the cost of the shortest path in $\mathcal{G}_{\text{start}}$ from $s_{0}$ to $s_{0}$ that visits all target-windows in $\Gamma$.
@@
At the beginning of Lazy BPRC, for every pair of segments $(\xi,\xi^{\prime})$ corresponding to different targets, we compute a lower bound on the cost of a feasible agent trajectory that intercepts $\xi$, then $\xi^{\prime}$, denoted as $c_{\text{seg}}(\xi,\xi^{\prime})$. To do so, we compute the shortest collision-free path in space from $\xi$ to $\xi^{\prime}$, ignoring time constraints, using the method from [16]. Let $c$ be the path’s distance traveled. Let $t$ be the start time of $\xi$, and let $t^{\prime}$ be the end time of $\xi^{\prime}$. We set $c_{\text{seg}}(\xi,\xi^{\prime})=c$ if $t+c/v_{\text{max}}\leq t^{\prime}$, and $c_{\text{seg}}(\xi,\xi^{\prime})=\infty$ otherwise.
@@
Next, given a tour $\Gamma$, we define a segment-graph $\mathcal{G}_{\text{seg}}=(\mathcal{V}_{\text{seg}},\mathcal{E}_{\text{seg}})$, where the set of nodes $\mathcal{V}_{\text{seg}}$ is the set of all segments whose target-windows are visited by $\Gamma$. For each edge $(\gamma,\gamma^{\prime})\in\mathcal{E}_{\text{tw}}$ traversed by $\Gamma$, we connect edges in $\mathcal{E}_{\text{seg}}$ from every segment of $\gamma$ to every segment of $\gamma^{\prime}$. The cost of each edge $(\xi,\xi^{\prime})$ is $c_{\text{seg}}(\xi,\xi^{\prime})$.
@@
For tours generated in the pricing problem (Section V-F), these $g$-values are computed as a byproduct of solving the pricing problem, and the lower bound $\underline{c}(\Gamma)$ is also strengthened using the minimum execution time for each prefix of $\Gamma$. For tours generated using the feasible solution generation method in Section V-H, we perform these $g$-value computations after generating the tours.
@@ V-F Pricing Problem
$\underline{c}_{\text{red}}(\Gamma)$ consists of the cost lower bound $\underline{c}(\Gamma)$, reduced by $\lambda_{i}$ for each target $i$ that $\Gamma$ visits. Large $\lambda_{i}$ can be interpreted as a bias toward tours that visit target $i$. Any tour that improves the optimal cost of the RMP must have negative reduced cost [11]. To find tours with negative reduced cost, we modify the labeling algorithm from [8] to handle obstacles.
@@ V-F2 Labeling Algorithm
  • $t^{\prime}={\textnormal{EFAT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}},t)$, where EFAT computes the earliest time at which a feasible agent trajectory could intercept $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time $t$. EFAT stands for “earliest feasible arrival time,” and we compute it using the method from [14].
@@
  • To compute each $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]$, we first check if $t^{\prime}>\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$; if so, $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]=\infty$, since an agent cannot arrive at $\xi_{i^{\prime},j^{\prime},k^{\prime}}$ by executing the partial tour represented by $l^{\prime}$. If $t^{\prime}\leq\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$, we compute
@@ V-G Computing Tour Cost via SPP-GCS
For a tour $\Gamma$, we compute its cost $c^{*}(\Gamma)$ via an SPP-GCS. At the beginning of Lazy BPRC, we decompose free space into convex regions $\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{n_{\text{reg}}}$, where $n_{\text{reg}}$ is the number of regions. We then define a GCS $\mathcal{G}_{\text{cs}}=(\mathcal{V}_{\text{cs}},\mathcal{E}_{\text{cs}})$, where the set of nodes $\mathcal{V}_{\text{cs}}$ consists of convex sets in space-time. For each region $\mathcal{A}$ in our free space decomposition, we have a region-node $\mathcal{X}_{\mathcal{A}}=\mathcal{A}\times\mathbb{R}$. For each target-window $\gamma_{i,j}$ visited by $\Gamma$, we have a window-node $\mathcal{X}_{i,j}$, consisting of the set of space-time points along $\tau_{i}$ within $\underline{t}_{i,j},\overline{t}_{i,j}]$. Since we assumed that a target’s velocity is constant within a time window, $\mathcal{X}_{i,j}$ is a line segment in space-time, which is a convex set. We refer to nodes in $\mathcal{G}_{\text{cs}}$ as GCS-nodes. An edge connects a set $\mathcal{X}$ to a set $\mathcal{X}^{\prime}$ if $\mathcal{X}$ intersects $\mathcal{X}^{\prime}$. We refer to a path in $\mathcal{G}_{\text{cs}}$ as a GCS-path. We say a GCS-path $P$ visits target-window $\gamma_{i,j}$ if $P$ contains $\mathcal{X}_{i,j}$.
@@
We search the GCS using a priority queue called OPEN, containing GCS-paths. We initialize OPEN with the GCS-path $(\mathcal{X}_{0,1})$ (i.e. a path that stays at the depot). Each GCS-path on OPEN has an $f$-value. The initial GCS-path $(\mathcal{X}_{0,1})$ has an $f$-value of 0, and we discuss the computation of $f$ for other GCS-paths later in the section. GCS-paths with smaller $f$-values have higher priority.
Each iteration pops a GCS-path $P$ from OPEN, then iterates over each GCS-node $\mathcal{X}$ adjacent to P[1]P[-1]. If $\mathcal{X}$ is a window-node $\mathcal{X}_{i,j}$, and $P$ has not visited the target-windows occurring before $\gamma_{i,j}$ in $\Gamma$, we discard $\mathcal{X}$. Furthermore, if $\mathcal{X}$ is a region-node, and $\mathcal{X}$ already occurs in $P$ after the final window-node in $P$, we discard $\mathcal{X}$. If we do not discard $\mathcal{X}$, we construct a successor GCS-path $P^{\prime}$ by appending $\mathcal{X}$ to $P$, then compute its $f$-value as follows.
@@ VII NUMERICAL RESULTS
We ran experiments on an Intel i9-9820X 3.3GHz CPU with 10 cores, hyperthreading disabled, and 128 GB RAM. We compared Lazy BPRC to two ablations. The first ablation, called “Non-Lazy BPRC,” is the same algorithm, but whenever we generate a label $l$ representing a tour $\Gamma$, and $l$ is not currently dominated, we set $\vec{g}_{\text{lb}}[1]=\vec{g}_{\text{ub}}[1]=c^{*}(\Gamma)$; if $\Gamma$ was unevaluated prior to this step, we evaluate $\Gamma$, update $\vec{g}_{\text{lb}}[1]$ and $\vec{g}_{\text{ub}}[1]$, then perform dominance checks again. The second ablation, called “No-Affine-Heuristic,” replaces our heuristic in the SPP-GCS associated with tour $\Gamma$ in Section V-G with the heuristic from FMC*. That is, when computing the $f$-value trajectory for a GCS-path $P^{\prime}$, the obstacle-unaware portion of the trajectory is required to intercept all target-windows in $\Gamma$ that are not visited by $P^{\prime}$, in sequence. Each algorithm parallelized successor generation in pricing and tour cost evaluation, the initial computation of pairwise distances between segments and segment-starts, and initial pairwise LFDT computations.
@@ VII-C Experiment 3: Varying the Capacity
We fixed $n_{\textnormal{tar}}$} to 9 and the map resolution to 30, then varied the capacity $d_{\text{max}}$ from 3 to 7. Fig. 3 (c) shows the results. Lazy BPRC shows smaller median runtime than both ablations, and also smaller min runtime than Non-Lazy BPRC. Lazy BPRC’s max runtime hits the time limit for $d_{\text{max}}\geq 4$. Note that No-Affine-Heuristic’s median runtime counterintuitively drops when we increase $d_{\text{max}}$ from 6 to 7. This occurs because in two instances, runtime became more than 5 times smaller when we increased $d_{\text{max}}$; runtime did not change significantly in the other 8 instances. The runtime dropped in these two instances for No-Affine-Heuristic because there were one or more tours whose evaluation required significant runtime for $d_{\text{max}}=6$, but simply never needed to be evaluated for $d_{\text{max}}=7$.

### 28 Mar 2026 v2

@@ I INTRODUCTION
Traditionally, the selection step within column generation (known as the restricted master problem) aims to minimize the sum of selected tours’ costs. In the MT-VRP-O, however, computing tour costs is expensive, since it requires collision-free motion planning. Our key idea is to instead perform column generation using cheap-to-compute lower bounds on tour costs. We compute these lower bounds by solving a motion planning problem with relaxed continuity constraints. Thus, we incorporate an outer alternation between (i) column generation using lower bounds on tour costs, and (ii) lazily evaluating only the costs of tours selected by column generation. We evaluate the cost of a tour by searching for a shortest path in a Graph of Convex Sets (GCS) [12]; we use our continuity relaxation strategy to provide a heuristic for the search. If column generation selects a set of tours whose costs have all been evaluated, we terminate the alternation between (i) and (ii). Our numerical results show that Lazy BPRC runs up to 46 times faster than a non-lazy ablation, and up to 26 times faster than an ablation using an existing obstacle-unaware heuristic [13].
@@ V-D Computing Lower Bound on Tour Cost
This section describes how we compute the lower bound $\underline{c}(\Gamma)$ for a tour $\Gamma$ in Section V. Our method, illustrated in Fig. 2, extends the procedure from BPRC [8] to handle obstacles. At the beginning of Lazy BPRC, we divide each target-window $\gamma_{i,j}$ into segments, where $\xi_{i,j,k}=(i,[\underline{t}_{i,j,k},\overline{t}_{i,j,k}])$ denotes the $k$th segment of $\gamma_{i,j}$, and ${\textnormal{Segments}}(\gamma_{i,j})$ is the set of segments of $\gamma_{i,j}$. To determine the number of segments per target-window, we first specify a number of segments to allocate per target, denoted as $n_{\text{seg,tar}}$. Then for each target $i$, we allocate segments to its windows using the formula from BPRC [8], which gives more segments to longer windows. The depot gets a single segment $\xi_{0}$. For a target-window $\gamma$, $n_{\text{seg}}(\gamma)$ is the number of segments allocated to $\gamma$. The segment indices for a target-window are ordered in increasing order of start time. An agent trajectory $\tau_{\textnormal{a}}$} intercepts segment $\xi_{i,j,k}$ if $\tau_{\textnormal{a}}$} intercepts target $i$ at a time $t\in[\underline{t}_{i,j,k},\overline{t}_{i,j,k}]$.
@@
Figure 2: Computing bounds on the cost of an example tour $\Gamma$. To compute the lower bound $\underline{c}(\Gamma)$, we divide each target-window visited by $\Gamma$ into segments. We then construct a segment-graph $\mathcal{G}_{\text{seg}}$, where the nodes are the segments, and an edge connects every segment of $\Gamma[n]$ to every segment of $\Gamma[n+1]$. The edge cost from segment $\xi$ to $\xi^{\prime}$ is the distance traveled along the shortest path in space from $\xi$ to $\xi^{\prime}$, if this path satisfies the relaxed timing constraints from in Section V-D, and $\infty$ otherwise. $\underline{c}(\Gamma)$ is the cost of the shortest path in $\mathcal{G}_{\text{seg}}$ from $\xi_{0}$ to $\xi_{0}$ visiting all target-windows in $\Gamma$. To compute the upper bound $\overline{c}$, we construct a segment-start-graph $\mathcal{G}_{\text{start}}$, where the nodes are the starting points of the segments, and an edge connects every segment-start of $\Gamma[n]$ to every segment-start of $\Gamma[n+1]$. The edge cost from $s$ to $s^{\prime}$ is distance traveled by a feasible minimum-distance agent trajectory from $s$ to $s^{\prime}$, if such a trajectory exists or if $s^{\prime}=s_{0}$, and $\infty$ otherwise. Our upper bound is the cost of the shortest path in $\mathcal{G}_{\text{start}}$ from $s_{0}$ to $s_{0}$ that visits all target-windows in $\Gamma$.
@@
For every pair of segments $(\xi,\xi^{\prime})$ corresponding to different targets, we compute a lower bound $c_{\text{seg}}(\xi,\xi^{\prime})$ on the cost of a feasible agent trajectory that intercepts $\xi$, then $\xi^{\prime}$. To do so, we compute the shortest collision-free path in space from $\xi$ to $\xi^{\prime}$, ignoring time constraints, via the method from [16]. Let $c$ be the path’s distance traveled. Let $t$ be the start time of $\xi$, and let $t^{\prime}$ be the end time of $\xi^{\prime}$. We set $c_{\text{seg}}(\xi,\xi^{\prime})=c$ if $t+c/v_{\text{max}}\leq t^{\prime}$, and $c_{\text{seg}}(\xi,\xi^{\prime})=\infty$ otherwise.
@@
Next, given a tour $\Gamma$, we define a segment-graph $\mathcal{G}_{\text{seg}}=(\mathcal{V}_{\text{seg}},\mathcal{E}_{\text{seg}})$, where the set of nodes $\mathcal{V}_{\text{seg}}$ is the set of all segments whose target-windows are visited by $\Gamma$. For each edge $(\Gamma[n],\Gamma[n+1])\in\mathcal{E}_{\text{tw}}$ traversed by $\Gamma$, we connect edges in $\mathcal{E}_{\text{seg}}$ from every segment of $\Gamma[n]$ to every segment of $\Gamma[n+1]$. The cost of each edge $(\xi,\xi^{\prime})$ is $c_{\text{seg}}(\xi,\xi^{\prime})$.
@@
For tours generated in the pricing problem (Section V-F), these $g$-values are computed as a byproduct of solving the pricing problem. For tours generated using the feasible solution generation method in Section V-H, we perform these $g$-value computations after generating the tours.
@@ V-F Pricing Problem
$\underline{c}_{\text{red}}(\Gamma)$ consists of the cost lower bound $\underline{c}(\Gamma)$, reduced by $\lambda_{i}$ for each target $i$ that $\Gamma$ visits. Large $\lambda_{i}$ can be interpreted as a bias toward tours that visit target $i$. To improve the RMP’s optimal cost, $\mathcal{F}_{\text{price}}$ must contain a tour with negative reduced cost [11]. To find tours with negative reduced cost, we modify the labeling algorithm from [8] to handle obstacles.
@@ V-F2 Labeling Algorithm
  • $t^{\prime}={\textnormal{EFAT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}},t)$, where EFAT is the earliest time at which a feasible agent trajectory can intercept $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time $t$. EFAT stands for “earliest feasible arrival time,” and we compute it using the method from [14].
@@ V-G Computing Tour Cost via SPP-GCS
At the beginning of Lazy BPRC, we decompose free space into convex regions $\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{n_{\text{reg}}}$, where $n_{\text{reg}}$ is the number of regions. We then define a GCS $\mathcal{G}_{\text{cs}}=(\mathcal{V}_{\text{cs}},\mathcal{E}_{\text{cs}})$, where the set of nodes $\mathcal{V}_{\text{cs}}$ consists of convex sets in space-time. For each region $\mathcal{A}$ in our free space decomposition, we have a region-node $\mathcal{X}_{\mathcal{A}}=\mathcal{A}\times\mathbb{R}$. For each target-window $\gamma_{i,j}$ visited by $\Gamma$, we have a window-node $\mathcal{X}_{i,j}$, consisting of the set of space-time points along $\tau_{i}$ within $\underline{t}_{i,j},\overline{t}_{i,j}]$. Since we assumed that a target’s velocity is constant within a time window, $\mathcal{X}_{i,j}$ is a line segment in space-time, which is a convex set. We refer to nodes in $\mathcal{G}_{\text{cs}}$ as GCS-nodes. An edge connects a set $\mathcal{X}$ to a set $\mathcal{X}^{\prime}$ if $\mathcal{X}$ intersects $\mathcal{X}^{\prime}$. We refer to a path in $\mathcal{G}_{\text{cs}}$ as a GCS-path. We say a GCS-path $P$ visits target-window $\gamma_{i,j}$ if $P$ contains $\mathcal{X}_{i,j}$.
@@
We search the GCS using a priority queue called OPEN, containing GCS-paths. We initialize OPEN with the GCS-path $(\mathcal{X}_{0,1})$, i.e. a path that stays at the depot. Each GCS-path on OPEN has an $f$-value. The initial GCS-path $(\mathcal{X}_{0,1})$ has an $f$-value of 0; we discuss the computation of $f$ for other GCS-paths shortly. GCS-paths with smaller $f$-values have higher priority.
Each iteration pops a GCS-path $P$ from OPEN, then iterates over each GCS-node $\mathcal{X}$ adjacent to P[1]P[-1]. If $\mathcal{X}$ is a window-node $\mathcal{X}_{i,j}$, and $P$ has not visited the target-windows occurring before $\gamma_{i,j}$ in $\Gamma$, we discard $\mathcal{X}$. If $\mathcal{X}$ is a region-node, and $\mathcal{X}$ already occurs in $P$ after the final window-node in $P$, we discard $\mathcal{X}$. If we do not discard $\mathcal{X}$, we construct a successor GCS-path $P^{\prime}$ by appending $\mathcal{X}$ to $P$, then compute its $f$-value as follows.
@@ VII NUMERICAL RESULTS
We ran experiments on an Intel i9-9820X 3.3GHz CPU with 10 cores, hyperthreading disabled, and 128 GB RAM. We compared Lazy BPRC to two ablations. The first ablation, called “Non-Lazy BPRC,” is the same algorithm, but whenever we generate a label $l$ representing a tour $\Gamma$, and $l$ is not currently dominated, we set $\vec{g}_{\text{lb}}[1]=\vec{g}_{\text{ub}}[1]=c^{*}(\Gamma)$; if $\Gamma$ was unevaluated prior to this step, we evaluate $\Gamma$, update $\vec{g}_{\text{lb}}[1]$ and $\vec{g}_{\text{ub}}[1]$, then perform dominance checks again. The second ablation, called “No-Affine-Heuristic,” replaces our heuristic in the SPP-GCS associated with tour $\Gamma$ in Section V-G with the heuristic from FMC*. That is, when computing the $f$-value trajectory for a GCS-path $P^{\prime}$, the obstacle-unaware portion of the trajectory is required to intercept all target-windows in $\Gamma$ unvisited by $P^{\prime}$, in sequence. Each algorithm parallelized successor generation in pricing and tour cost evaluation, the initial computation of pairwise distances between segments and segment-starts, and initial pairwise LFDT computations.
@@ VII-C Experiment 3: Varying the Capacity
We fixed $n_{\textnormal{tar}}$} to 9 and the map resolution to 30, then varied the capacity $d_{\text{max}}$ from 3 to 7. Fig. 3 (c) shows the results. Lazy BPRC shows smaller median runtime than both ablations, and also smaller min runtime than Non-Lazy BPRC. Lazy BPRC’s max runtime hits the time limit for $d_{\text{max}}\geq 4$.
Note that No-Affine-Heuristic’s median runtime counterintuitively drops when we increase $d_{\text{max}}$ from 6 to 7. This occurs because in two instances, runtime became more than 2 times smaller when we increased $d_{\text{max}}$; runtime did not change as significantly in the other 8 instances. The runtime dropped in these two instances for No-Affine-Heuristic because there were one or more tours whose evaluation required significant runtime for $d_{\text{max}}=6$, but simply never needed to be evaluated for $d_{\text{max}}=7$.